OpenStudy (anonymous):

Solve the equation, using the domain of x = {-1, 0, 2}. 2x + 3y = 7

OpenStudy (anonymous):

You simply plug in the allowed x values and get the corresponding y-values. Ex: when x=-1 you have 2(-1) +3y=7 -2+3y=7 3y=9 y=3 so you have the solution (-1, 3) When x=0 2(0)+3y=7 3y=7 y=7/3 So you have the solution (0, 7/3) When x=2 2(2)+3y=7 4+3y=7 3y=3 y=1 So you have the solution (2, 1) Or another way to write this is y={3, 7/3, 1} The fancy brackets { } denote a set of numbers So y={3, 7/3, 1} is your range (range is the allowed values of y)

OpenStudy (anonymous):

Or, solve the equation for y first: \[2x + 3y = 7 \rightarrow y=\frac{7-2x}{3}.\] Then input your domain: \[y=\frac{7-2\left \{-1, 0, 2 \right \}}{3}\] and evaluate.